OpenStudy (anonymous):

What is 8(1+5x)+5=13+5x

OpenStudy (anonymous):

8(1+5x)+5=13+5x 8+40x+5=13+5x 40x+13=13+5x -13 -13 40x=5x -5x -5x 35x=0 so x=0

OpenStudy (uri):